OpenStudy (anonymous):

5. What is a cubic polynomial function in standard form with zeros 1, –2, and 2? (1 point) = x3 + x2 – 3x + 4 = x3 + x2 – 4x – 2 = x3 + x2 + 4x + 4 = x3 – x2 – 4x + 4

OpenStudy (anonymous):

Cubic stands for ^3

OpenStudy (anonymous):

Now, just factor each of them to find the zeros.

OpenStudy (anonymous):

is the ansewer c

OpenStudy (anonymous):

Just wait, let me check.

OpenStudy (anonymous):

you can always use http://www.wolframalpha.com/ to help you.

jimthompson5910 (jim_thompson5910):

y = x^3+x^2+4x+4 y = (1)^3+(1)^2+4(1)+4 ... plug in x = 1 y = 1 + 1 + 4(1) + 4 y = 1 + 1 + 4 + 4 y = 10 Since this is NOT zero, this means x = 1 is NOT a root or zero

jimthompson5910 (jim_thompson5910):

so it can't be C

OpenStudy (anonymous):

Oh, that's a simpler way to do it when you have the zeros given. Didn't see that :o

OpenStudy (anonymous):

i need help i dont get it

OpenStudy (anonymous):

The zeros are where the equation crosses on the x-axis on the graph.

OpenStudy (anonymous):

The normal way (or the way I learnt it) is you have to factor your equation first so then you can find the zeros but @jim_thompson5910 's idea is way simpler.

jimthompson5910 (jim_thompson5910):

if x = 1 is a zero, then plugging it into some equation *should* give you y = 0

jimthompson5910 (jim_thompson5910):

ex: if y = x-1 then x = 1 is a zero because y = x-1 y = 1-1 .. plug in x = 1 y = 0 So this shows that x = 1 is a root or zero of y = x-1

OpenStudy (anonymous):

would it be b then

jimthompson5910 (jim_thompson5910):

what do you get when you plug x = 1 into y = x^3+x^2-4x-2

OpenStudy (anonymous):

help pleezzz

jimthompson5910 (jim_thompson5910):

y = x^3+x^2-4x-2 y = (1)^3+(1)^2-4(1)-2 y = ???

OpenStudy (anonymous):

y= -8

jimthompson5910 (jim_thompson5910):

that is NOT y = 0, so choice B is also out

jimthompson5910 (jim_thompson5910):

btw it should be y = -4, but that's still not 0

OpenStudy (anonymous):

so what is the answer

jimthompson5910 (jim_thompson5910):

well you've eliminated B and C so far

jimthompson5910 (jim_thompson5910):

try out A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it a

jimthompson5910 (jim_thompson5910):

what do you get when you plug in x = 1

OpenStudy (anonymous):

i got 3

jimthompson5910 (jim_thompson5910):

which is not 0, so x = 1 is not a root for choice A

OpenStudy (anonymous):

so its d

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np